• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 17:52:56 on July24,2006

    Entry number 175144

    keyword=heat leak fixed and temp sensors calibrated

    During the day access, Mike Seely from the cryotarget group fixed a heat leak in the switch panel. He also did calibration of the two new CERNEX temperature sensors in loop 3 (Heater_fore and Heater_aft). One now reads correct temperature and the other is a little over 1K too low.